# Audio-Synchronized Visual Animation

### v1

@@ 4.1 Preliminary: Text-to-Image Latent Diffusion
During inference, LDMs start from a random gaussian noise map ${\bm{z}}^{K}$, and iterate over $K$ reverse diffusion steps [17, 36] for cleaner latent ${\bm{z}}^{k-1}={\bm{z}}^{k}-\epsilon_{\theta}({\bm{z}}^{k},k,{\bm{\tau}})$, until ${\bm{z}}^{0}$ is obtained. LDMs then decode the latent into image ${\bm{x}}^{0}={\mathcal{D}}({\bm{z}}^{0})$ using the pre-trained decoder ${\mathcal{D}}(\cdot)$. For simplicity, we refer to the images by their latents ${\bm{z}}$ rather than ${\bm{x}}$ throughout the rest of the paper.
4.2 Proposed Architecture
In this work, we extend the capabilities of LDMs for ASVA, by focusing on learning video dynamics and temporal synchronization, unlike existing approaches [14, 37, 38] which primarily use audio to guide global semantics. We propose the Audio-Video Synchronized Diffusion model (AVSyncD), which builds upon a pretrained image LDM and integrates synchronized audio control and temporal layers for improved video consistency. The major component is a UNet $\epsilon_{\theta}$ aiming to denoise a video instead of image. The UNet is trained on synchronized audio-video pairs, with the first frame ${\bm{z}}_{1}$, the corresponding audio ${\bm{a}}$, and the CLIP encoded audio category name ${\bm{\tau}}$ as input conditions. The LDM denoising objective in Eq. 5 is applied to the remaining frames ${\bm{z}}_{2:rT}$ to be predicted. The architecture of AVSyncD is shown in Fig. 3 and discussed below.
Initial-frame Conditioning
@@ 4.2 Proposed Architecture
To condition LDM on input image, we feed initial frame’s latent ${\bm{z}}_{1}$ without noise into the UNet at every diffusion timestep $k$. For all subsequent frames, we adhere to the original LDM, using independently sampled noisy latents ${\bm{z}}_{2:rT}^{k}$ as inputs and predicting the added noise $\epsilon^{k}$.
Text Conditioning
@@
We retained the text cross attention layers in original LDMs [32] by conditioning on the audio category as a prompt. It however brings insignificant gains due to limited text diversity in training, see supplementary.
Audio Conditioning
@@
To facilitate audio-synchronized generation, we condition the generation on ImageBind audio embeddings [14]. ImageBind computes a audio classification token, ${\bm{a}}^{g}$, representing global semantics, and local patch tokens, ${\bm{a}}_{f,t}$, across $T_{a}$ timestamps. Original ImageBind only uses ${\bm{a}}^{g}$ for contrastive learning and discards ${\bm{a}}_{f,t}$. However, we found these frozen local tokens as efficient synchronization cues. We split the patch tokens temporally into $rT$ segments, corresponding to the same timestamps as the frames ${\bm{z}}_{1:rT}$ , and append the global token to each. Each frame ${\bm{z}}_{t}$ then learns both semantics and synchronization guidance from its audio segment ${\bm{a}}_{t}$ via cross-attention [41].
Spatial Convolution
@@
The original LDM’s pretrained spatial convolutional layers were frozen and used without modification.
First-frame Temporal Convolution
@@
Each spatial convolution block was augmented with a 1D temporal convolution layer (kernel size 3) to capture temporal dependencies. To better adhere to the starting image, ${\bm{z}}_{1}$, we adjusted the receptive field at frame $t$ to include frames $(1,t-1,t)$ as opposed to $(t-1,t,t+1)$. These temporal convolutions with first-frame lookup were applied to three components in the UNet, i.e., the input/output conv layers and all ResNet convs.
First-frame Spatial Attention
@@
We also leveraged the pretrained LDM’s spatial self-attentions. Following [20], we modified the frozen spatial attention layers to cross-attend to the first frame rather than self-attend to the current frame, by computing key-values from the first frame and queries from the current one.
Temporal Attention
@@
To effectively model long-range visual dependencies, we incorporated temporal attention layers [3]. Each frame index $t$ was converted into a sinusoidal positional embedding, added to the corresponding frame’s latents after a learnable projection [41]. Each frame’s local representation at spatial position ($h,w$), ${\bm{z}}_{hwt}$, was then updated by attending to all frames at the same position $({\bm{z}}_{hw1},{\bm{z}}_{hw2},\ldots,{\bm{z}}_{hw(rT)})$ through a standard attention mechanism.
[…]
Classifier-free Audio Guidance
@@
Classifier-free guidance [18] is a technique used to control the impact of input prompts on the generated outputs. We extended it to amplify audio guidance for improved synchronization. To accomplish this, we trained the model for both audio-conditioned and unconditioned generation, by randomly replacing ${\bm{a}}$ with a null audio embedding, ${\bm{a}}_{\emptyset}$, with a 20% likelihood. ${\bm{a}}_{\emptyset}$ was computed by encoding an all-zero audio melspectrogram via ImageBind. During inference, a factor $\eta\geq 1$ controls the audio guidance:
@@
Moreover, for comparison, we also trained an image-to-video variant of AVSyncD, named I2VD, by removing audio conditioning.
@@ 5 Experiments
We provide more experiment details in supplementary, including implementation details, experiment results, and sounded videos for comparison.
@@ 5.1 Implementation
Dataset
We experimented on 3 datasets. (1) AVSync15: Our high-quality dataset curated from VGGSound [7]. It is a class-balanced dataset, with 15 motion categories, 1350 training videos, and 150 test videos. We also assessed our data curation pipeline via experiments on uncurated VGGSS and automatically curated AVSync-AC. (2) Landscapes [21]: A conventional dataset for audio-to-video generation, with 9 environmental sound classes. We followed the split in  [33, 46] with 900 clips for training and 100 for testing. Since Landscapes is full of ambient sounds without synchronized video motion, we mainly use it to evaluate visual quality. (3) TheGreatestHits [27]: An audio-video dataset recording humans probing environments with a drumstick, with 733 videos for training and 244 for testing. The videos are clean, synchronized with audio at the moments of probing, but still contain a large portion of moments with static background. It also offers limited diversity, featuring a singular motion of probing.
Data Preprocess
We sampled 2-second synchronized audio-video pairs for experiments. Videos were sampled at 6 fps with 12 frames, and resized to 256$\times$256 on AVSync15/Landscapes or 128$\times$256 on TheGreatestHits. Following ImageBind [14], audios were sampled at 16kHz and converted into 128-d spectrograms.
Baselines
We first adopted a simple baseline, Static, by repeating the input frame into a video, then compared to several state-of-the-art audio-to-video methods. (1) Semantic audio-to-video generation (CoDi [38]): Image, text, and audio are encoded into a shared CLIP [30] space and summed, then cross attend to a video diffusion model trained on large-scale datasets. (2) Image-to-video generation (VideoCrafter [5]): A video diffusion model superior to CoDi on video generation but without audio inputs. It animates images by fusing CLIP-encoded image and text features into the model via modality-dependent cross attention layers. (3) Synchronized audio-to-video generation (TPoS [19], AADiff [22], TempoToken [46]): Audio is encoded into time-dependent segments and fused into frozen text-to-image [32] or text-to-video [42] models. We re-implemented AADiff, and used TPoS and TempoToken’s pretrained checkpoints on VGGSound and Landscapes. More details provided in supplementary.
Training & Evaluation
We adopted the pretrained Stable Diffusion-V1.5 [32] as diffusion model and ImageBind [14] as audio encoder. All models were primarily trained using Adam optimizer with a batch size of 64, a learning rate of 0.0001, with 37k/24k/16k iterations on AVSync15/Landscapes/TheGreatestHits, respectively. Evaluation metrics include FID, IA, IT, FVD, AlignSync, RelSync, and a user study, as detailed in Sec. 3.2. At test time, we evaluated on 3 videos uniformly sampled from each example. More details provided in supplementary.
@@ 5.2 Main Results
Figure 4: Qualitative results on three datasets. Sounded videos are provided in supplementary.
@@
Input Model FIDnormal-$\downarrow$ IAnormal-$\uparrow$ ITnormal-$\uparrow$ FVDnormal-$\downarrow$ AlignSyncnormal-$\uparrow$ RelSyncnormal-$\uparrow$ User Study
IQnormal-$\uparrow$ FCnormal-$\uparrow$ Syncnormal-$\uparrow$
@@
Input Model FIDnormal-$\downarrow$ IAnormal-$\uparrow$ ITnormal-$\uparrow$ FVDnormal-$\downarrow$ AignSyncnormal-$\uparrow$ RelSyncnormal-$\uparrow$
@@
Input Model FIDnormal-$\downarrow$ IAnormal-$\uparrow$ ITnormal-$\uparrow$ FVDnormal-$\downarrow$ AignSyncnormal-$\uparrow$ RelSyncnormal-$\uparrow$
@@
Dataset Comparison
In Tab. 1, the Static baseline is a good indicator of dataset attributes. It has similar IA/IT scores compared to Groundtruth, since Static is composed of a subset of groundtruth frames, which are obviously semantically aligned. FVD of Static on TheGreatestHits is relatively low, since the TheGreatestHits contains frequent static groundtruth clips without any moving objects. More importantly, the RelSync and AlignSync of Static gradually increase from AVSync15, to TheGreatestHits, to Landscapes, with those on Landscapes even surpassing the groundtruth. The fact that static videos perform well in terms of audio synchronization on TheGreatestHits and Landscapes datasets testify for the superiority of AVSync15 for audio-video synchronized generation.
Model Comparison
We compared to prior works on AVSync15 and Landscapes in Tabs. 1(a), 1(b), 3(a) and 3(b). In Tab. 1(a), CoDi achieves inferior results on almost all metrics. TPoS(I+T+A) shows strong image quality (FID), but is worse in video quality (FVD) and synchronization (RelSync). TempoToken, on the other hand, is better at synchronization rather than visual quality, likely due to lack of image input. AADiff is competitive on synchronization but extremely bad on image quality (FID). This is expected as AADiff adjusts each frame using audio amplitude, producing visual changes that highly correlate to audio changes temporally but may be overwhelmed by noises, as shown in Fig. 3(a). On Landscapes, its similar FVD to Static but abnormally higher AlignSync and RelSync also suggest that it only applied minor modifications to the input image due to lack of sound changes, as shown in Fig. 3(b). Without audio input, VideoCrafter performs poorly on synchronization. It also has difficulty faithfully adhering to the input frame, as in Fig. 3(a). Our AVSyncD, however, achieves the best animation results on almost all metrics. On Landscapes, AVSyncD also performs the best on FID and FVD, with other metrics being similar to the groundtruth. We provide qualitative comparison in Fig. 4.
User Study
We invited 15 participants to compare 4 animation models with top overall performance (VideoCrafter, AADiff, I2VD, AVSyncD) on AVSync15, based on 3 metrics in Tab. 1. The 4 models generated videos conditioned on the same test examples (audio+image). Each test example was independently evaluated by 2 participants to select their most preferred generation (vote) on each metric. In total, we evaluated all 150 test examples on AVSync15, collecting 150$\times$2=300 votes on each metric. Tab. 1(a) shows votes each model received. AVSyncD received the most votes on all metrics.
@@ 5.3 Ablation Studies
Audio Conditioning
In Tab. 1, AVSyncD almost always outperforms I2VD on all metrics, expecially AlignSync and RelSync. On Landscapes, however, AVSyncD does not improve RelSync, further suggesting lack of synchronization cues on Landscapes. These results validates AVSyncD can effectively utilize audio as guidance to improve visual quality as well as synchronization.
Audio Guidance
Tab. 1 shows increasing the audio guidance factor $\eta$ from 1 to 4 improves FID, IA, and FVD on all three datasets. As expected, audio guidance also improved AlignSync and RelSync significantly on AVSync15 and TheGreatestHits, but not on the less synchronized Landscapes dataset. Prior works [37, 22, 19] claimed that increasing audio amplitude can also lead to stronger visual effects. We compare this approach with our audio guidance in Fig. 4(a). Audio guidance offers a better control mechanism than audio amplitude.
@@
FF-Conv FF-Attn FIDnormal-$\downarrow$ FVDnormal-$\downarrow$ AignSyncnormal-$\uparrow$
@@
Dataset AC MC FIDnormal-$\downarrow$ FVDnormal-$\downarrow$ AignSyncnormal-$\uparrow$
@@
Effect of First-frame (FF) Lookups
We validate the effectiveness of First-frame (FF) Lookups by replacing them with the standard temporal convolutions or spatial self-attentions. Tab. 2(a) shows their effectiveness.
Effect of Data Curation
We trained AVSyncD on randomly sampled subsets from VGGSS and AVSync-AC, ensuring equal training data scale and balanced category distributions. Results in Tab. 2(b) justified our data curation pipeline.
5.4 Applications and Extensions
Despite AVSync15 as a relatively small dataset, we found AVSyncD trained on it can generalize well to many applications. See more results in supplementary.
@@ 5.4 Applications and Extensions
(b) Animate target objects on internet images with audios. Cartridges ejection effects when firing a gun are better viewed in videos in supplementary.
@@
Animate Generated Images
When lacking an image as input, we can use existing image generators to generate the image, which AVSyncD can also animate. Fig. 4(b) shows animations on images generated by StableDiffusion-V1.5 [32].
Animate Contents from Internet
AVSyncD can also generalize well to unseen images and audios from internet, as shown in Fig. 4(c).
Control Animations with (Un)related Audios
We can control the motion of an image to follow desired audios, e.g., animate a baby to not only cry, but also bark like a dog or roar like a lion, as seen in Fig. 5(a). When there is no object related to the audio, the animations do not demonstrate corresponding motion.
Animate Target Objects with Audios
When multiple objects of interests exist in the image, a scenario not existing in training data, we can still use audios to only animate the related target object, as shown in Fig. 5(b).

### v2

@@ 4.1 Preliminary: Text-to-Image Latent Diffusion
${\mathcal{L}}_{\textit{LDM}}={\mathbb{E}}_{{\bm{z}},\epsilon\sim{\mathcal{N}}(0,1),k}\left[\|\epsilon-\epsilon_{\theta}({\bm{z}}^{k},k,{\bm{\tau}})\|_{2}^{2}\right]$ (5)
During inference, LDMs start from a random Gaussian noise map ${\bm{z}}^{K}$, and iterate over $K$ reverse diffusion steps [17, 38] to denoise the latents ${\bm{z}}^{k-1}={\bm{z}}^{k}-\epsilon_{\theta}({\bm{z}}^{k},k,{\bm{\tau}})$, until ${\bm{z}}^{0}$ is obtained. LDMs then decode the latent into an image ${\bm{x}}^{0}={\mathcal{D}}({\bm{z}}^{0})$ using the pre-trained decoder ${\mathcal{D}}(\cdot)$. For simplicity, we refer to the images by their latents ${\bm{z}}$ rather than ${\bm{x}}$ throughout the rest of the paper.
4.2 Proposed Architecture
In this work, we extend the capabilities of LDMs for ASVA, by focusing on learning video dynamics and temporal synchronization, unlike existing approaches [14, 39, 40] which primarily use audio to control global semantics. We propose the Audio-Video Synchronized Diffusion model (AVSyncD), which builds upon a pre-trained image LDM and integrates synchronized audio control and temporal layers for improved video consistency. The major component is a UNet $\epsilon_{\theta}$ aiming to denoise a video instead of an image. The UNet is trained on synchronized audio-video pairs, with the first frame ${\bm{z}}_{1}$, the corresponding audio ${\bm{a}}$, and the CLIP encoded audio category name ${\bm{\tau}}$ as input conditions. The LDM denoising objective in Eq. 5 is applied to the remaining frames ${\bm{z}}_{2:rT}$ to be predicted. The architecture of AVSyncD is shown in Fig. 3 and discussed below. Further details, for example, highlighting the different attention modules used in the architecture are described in Suppl. Sec. 3.2.
@@ 4.2 Proposed Architecture
Initial-frame Conditioning To condition LDM on an input image, we feed its latent ${\bm{z}}_{1}$ without noise into the UNet at every diffusion timestep $k$. For all subsequent frames, we adhere to the original LDM, using independently sampled noisy latents ${\bm{z}}_{2:rT}^{k}$ as inputs and predicting the added noise $\epsilon^{k}$.
@@
Text Conditioning We retain the text cross-attention layers in the original LDMs [34] (without finetuning) to condition the model on the audio category. However, due to the limited text diversity in training, class conditioning does not bring significant gains (see Suppl. Sec. 6.5).
@@
Audio Conditioning To facilitate audio-synchronized generation, we condition the generation on ImageBind audio embeddings [14]. ImageBind computes an audio classification token, ${\bm{a}}^{g}$, representing global semantics, and local patch tokens, ${\bm{a}}_{f,t}$, across $T_{a}$ timestamps. The original ImageBind only uses ${\bm{a}}^{g}$ for contrastive learning and discards ${\bm{a}}_{f,t}$. However, we found these frozen local tokens as efficient synchronization cues. We split the patch tokens temporally into $rT$ segments, corresponding to the same timestamps as the frames ${\bm{z}}_{1:rT}$, and append the global token to each. Each frame ${\bm{z}}_{t}$ then learns both semantics and synchronization guidance from its audio segment ${\bm{a}}_{t}$ via cross-attention [43]. In Suppl. Sec. 6.7, we compared ImageBind with CLAP and BEATs as encoders.
@@
Spatial Convolution The original LDM’s pre-trained spatial convolutional layers were frozen and used without modification.
@@
First-frame Temporal Convolution Each spatial convolution block was augmented with a 1D temporal convolution layer (kernel size 3) to capture temporal dependencies. To better adhere to the starting image, ${\bm{z}}_{1}$, we adjusted the receptive field at frame $t$ to include frames $(1,t-1,t)$ as opposed to $(t-1,t,t+1)$. These temporal convolutions with first-frame lookup were applied to three components in the UNet, namely the input/output conv layers and all ResNet convs.
@@
First-frame Spatial Attention We also leveraged the pre-trained LDM’s spatial self-attentions. Following [21], we modified the frozen spatial attention layers to cross-attend to the first frame rather than self-attend to the current frame, by computing the key-value pairs from the first frame and the queries from the current one.
@@
Temporal Attention To effectively model long-range visual dependencies, we incorporated temporal attention layers [2]. Each frame index $t$ was converted into a sinusoidal positional embedding, added to the corresponding frame’s latents after a learnable projection [43]. Each frame’s local representation at spatial position ($h,w$), ${\bm{z}}_{hwt}$, was then updated by attending to all frames at the same position $({\bm{z}}_{hw1},{\bm{z}}_{hw2},\ldots,{\bm{z}}_{hw(rT)})$ through a standard attention mechanism.
@@
Classifier-free Audio Guidance Classifier-free guidance [18] is a technique used to control the impact of input prompts on the generated outputs. We extended it to amplify audio guidance for improved synchronization. To accomplish this, we trained the model for both audio-conditioned and unconditioned generation, by randomly replacing ${\bm{a}}$ with a null audio embedding, ${\bm{a}}_{\emptyset}$, with a 20% likelihood. ${\bm{a}}_{\emptyset}$ was computed by encoding an all-zero audio mel spectrogram via ImageBind. During inference, a factor $\eta\geq 1$ controls the audio guidance:
@@ 5.1 Implementation
Dataset We conducted experiments on three datasets. AVSync15: Our high-quality dataset curated from VGGSound [7], with 15 balanced categories, 1350 training videos, and 150 test videos. We also assessed our data curation pipeline by comparing it to models trained on Landscapes, TheGreatestHits, VGGSS and our AVSync-AC. Landscapes [22] is composed of 9 environmental sound classes. We followed the split in [35, 48] with 900 clips for training and 100 for testing. Since Landscapes is full of ambient sounds without synchronized video motion, we mainly use it to evaluate visual quality. TheGreatestHits [29] is an audio-video dataset recording humans probing environments with a drumstick, with 733 videos for training and 244 for testing. The videos are synchronized with audio at the moments of impact but contain lots of static moments. It also offers limited diversity, featuring a singular motion of probing. VGGSS and our AVSync-AC were described in Sec. 3.
Data Preprocess We sampled 2-second synchronized audio-video pairs for experiments. Videos were sampled at 6 fps with 12 frames, and resized to 256$\times$256 on AVSync15/Landscapes or 128$\times$256 on TheGreatestHits. Following ImageBind [14], audios were sampled at 16kHz and converted into 128-d spectrograms.
Baselines We first adopted a simple Static baseline by repeating the input frame into a video, then compared it to several state-of-the-art works. (1) Semantic audio-to-video generation (CoDi [40]): Image, text, and audio are encoded into a shared CLIP [32] space and summed, then fused into a video diffusion model trained on large-scale datasets. (2) Image-to-video generation (VideoCrafter [5]): A superior video diffusion model however without audio inputs. It animates images by fusing CLIP-encoded image and text features into the model via modality-dependent cross-attention layers. (3) Synchronized audio-to-video generation (TPoS [20], AADiff [23], TempoToken [48]): Audio is encoded into time-dependent segments and fused into frozen text-to-image [34] or text-to-video [44] models. We re-implemented AADiff, and used TPoS and TempoToken’s pretrained checkpoints on VGGSound and Landscapes. More details are provided in Suppl. Sec. 4.
Training & Evaluation We adopted the pretrained Stable Diffusion-V1.5 [34] as the diffusion model and ImageBind [14] as the audio encoder. All models were trained using Adam optimizer with a batch size of 64 and a learning rate of 0.0001. Besides metrics in Sec. 3.2, we also provide results using metric in [26] in Suppl. Sec. 3.4. We evaluated on 3 clips uniformly sampled from each video.
@@ 5.2 Main Results
Figure 4: Qualitative results on three datasets.
@@
[…]
@@
[…]
@@
[…]
@@
Dataset Comparison In Tab. 1(c), the Static baseline is a good indicator of dataset attributes. It has similar IA/IT scores compared to Groundtruth, since Static is composed of a subset of groundtruth frames, which are obviously semantically aligned. FVD of Static on TheGreatestHits is relatively low, since the TheGreatestHits contains frequent static ground truth clips without any moving objects. More importantly, the RelSync and AlignSync of Static gradually increase from AVSync15, to TheGreatestHits, to Landscapes, with those on Landscapes even surpassing the ground truth. The fact that static videos perform well in terms of audio synchronization on TheGreatestHits and Landscapes datasets testify to the superiority of AVSync15 for audio-video synchronized generation.
Model Comparison We compared to prior works on AVSync15 and Landscapes in Tabs. 1(a), 1(b), 4(a) and 4(b). In Tab. 1(a), CoDi achieves inferior results on almost all metrics. TPoS(I+T+A) shows strong image quality (FID), but is worse in video quality (FVD) and synchronization (RelSync). TempoToken, on the other hand, is better at synchronization rather than visual quality, likely due to the lack of image input. AADiff is competitive on synchronization but extremely bad on image quality (FID). This is expected as AADiff adjusts each frame using audio amplitude, producing visual changes that highly correlate to audio changes temporally but may be overwhelmed by noises, as shown in Fig. 4(a). On Landscapes, its similar FVD to Static but abnormally higher AlignSync and RelSync also suggest that it only applied minor modifications to the input image due to lack of sound changes, as shown in Fig. 4(b). Without audio input, VideoCrafter performs poorly on synchronization. It also has difficulty faithfully adhering to the input frame, as in Fig. 4(a). AVSyncD achieves the best animation results on almost all metrics. On Landscapes, AVSyncD also performs the best on FID and FVD, with other scores being similar to ground truth.
User Study We invited 15 participants to compare 4 animation models with top overall performance (VideoCrafter, AADiff, I2VD, AVSyncD) on AVSync15, based on 3 metrics in Tab. 1(c). The 4 models generated videos conditioned on the same test examples (audio+image). Each test example was independently evaluated by 2 participants to select their most preferred generation (vote) on each metric. In total, we evaluated all 150 test examples on AVSync15, collecting 150$\times$2=300 votes on each metric. Tab. 1(a) shows votes each model received.
@@ 5.3 Ablation Studies
Audio Conditioning In Tab. 1(c), AVSyncD outperforms I2VD, especially on AlignSync and RelSync. AVSyncD did not improve RelSync on Landscapes probably due to the lack of synchronization cues on the dataset itself. These results show that audio condition enhances generation quality and synchronization.
Audio Guidance Tab. 1(c) shows increasing the audio guidance factor $\eta$ from 1 to 4 improves FID, IA, and FVD on all three datasets. As expected, audio guidance also improved AlignSync and RelSync significantly on AVSync15 and TheGreatestHits, but not on the less synchronized Landscapes dataset. Prior works [39, 23, 20] claimed that increasing audio amplitude can also lead to stronger visual effects. We compare this approach with our audio guidance in Fig. 5(a). Audio guidance offers a better control mechanism than audio amplitude.
First-frame (FF) Lookups We validated FF Lookups by replacing them with standard temporal convolutions or spatial self-attention in Tab. 2(a).
Data Curation We compared to AVSyncD trained on random subsets from VGGSS and AVSync-AC with equal data scale and balanced categories in Tab. 2(b).
5.4 Applications and Extensions
@@
[…]
@@
[…]
@@
(b) Animate target objects on internet images with audios.
@@
Animate Generated Images When lacking an image as input, we can use existing image generators to generate the image, which AVSyncD can also animate. Fig. 5(b) shows animations on images generated by StableDiffusion-V1.5 [34].
Animate Contents from Internet AVSyncD can also generalize well to unseen images and audio, as shown in Fig. 5(c).
Control Animations with (Un)Related Audios We can control the motion of an image to follow desired audio, e.g., animate a baby to not only cry but also bark like a dog or roar like a lion, as seen in Fig. 6(a). When there is no object related to the audio, the animations do not demonstrate corresponding motion.
Animate Target Objects with Audios When multiple objects exist in the image, a scenario not existing in training data, we can still use audios to only animate the related target object, as shown in Fig. 6(b).